OpenStudy (anonymous):

find the limit of lim x->pie/2 [2 sin(pie+cotx/cosx+2sinx)]

OpenStudy (anonymous):

\[x \rightarrow 0.5 \pi(2\sin(\pi+\frac{cosx}{sinx cosx}+2sinx))\] \[x \rightarrow 0.5 \pi(2sin(\pi+\frac{1}{sinx}+2sinx)) \] \[x \rightarrow \pi/2 (\sin(x))=1\] \[x \rightarrow 0.5 \pi(2sin(\pi+2\frac{1}{2})) \] Just calculate it